OpenStudy (anonymous):

write the expression logbsqrt(128/27) in terms of A and C when A=logb2 and C=logb3

OpenStudy (anonymous):

use the fact that \[128=2^7\] and \[27=3^3\] to rewrite \[\log(\sqrt{\frac{2^7}{3^3})}\]

OpenStudy (anonymous):

then use \[\sqrt{x}=x^{\frac{1}{2}}\] to rewrite it as \[\log(\frac{2^7}{3^2})^{\frac{1}{2}}\]

OpenStudy (anonymous):

then \[\frac{1}{2}(\log(2^7)-\log(3^3))\]

OpenStudy (anonymous):

then \[\frac{1}{2}(7\log(2)-3\log(3))\] and finally you have what you want, an expression in \[\log(2)\text { and } \log(3)\]

OpenStudy (anonymous):

so your "final answer" is \[\frac{1}{2}(7A-3C)\]

OpenStudy (anonymous):

hello fred!

OpenStudy (anonymous):

when do i get to level 101?

OpenStudy (anonymous):

hahaha ur pretty amazing. Always help me.

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

yw, hope steps are clear